Please with steps:


Consider the mixing of 5.1 L of CO2 (g) and 3.1 L of H2O (g) at 32°C and 850 mm Hg. Determine the mass of O2 (g) that can be produced from the unbalanced reaction:


CO2 (g) + H2O (g) = C4H10 (l) + O2 (g)


Mass=...............g





1
Expert's answer
2021-03-09T05:00:01-0500

CO2 (g) + H2O (g) = C4H10 (l) + O2 (g)

Total Volume = 5.1L + 3.1L = 8.2L

n = PV/RT = 850 mm Hg x 5.1 / 8.314 J/mol x 302k

= 1.73mol

O2 = 32g/mol

mass = mole X molar mass

= 1.73mol x 32g/mol

= 55.36g